Assume that the banking system is loaned up and that any open-market purchase by the Fed directly increases reserves in the banks. If the required reserve ratio is 0.2, by how much could the money supply expand if the Fed purchased $2 billion worth of bonds?
1
Expert's answer
2016-05-23T10:03:03-0400
If the required reserve ratio is rr = 0.2, and if the Fed purchased $2 billion worth of bonds, then the money multiplier in this case is mm = 1/rr = 1/0.2 = 5, so the money supply could expand by 5*$2 billion = $10 billion.